a crickter can throw aball to maximum horzontal distance  of 180m . calculate the max. vertical height to which he can  throw the ball 9g=10m\s2)

Grade:12

1 Answers

Ritika Bagokar
7 years ago
Rmax=180    therefore angle of projection should be 45°
u^2/g=180  
u^2=1800 
u=30√2
 
for maximum height angle of projection must be 90°  
Hmax=u^2/2g  =30√2*30√2/20  =30m  
hence maximum height to which he can throw a ball is 30m